A) 1.92 diopters B) 2.75 diopters C) 0.4 diopters D) 0.96 diopters E) 5.052632 diopters -80 cm, what would be his prescriptions lens power diopters 2. A reaistor is connected what factor does the power dissipated in the circuit change to a battery. If you replace the resistor with one t A) 4 B) 8 C) 2 D) 1/2 E)14 change? 3. Three resiators connected in parallel have individual values ot 4.0, 6.0 and 10.0 G respectively. It this combination ia what is the current in the 2 a resistor? connected in series with a 12.0 V battery 2.09 4.02 6.0 ? 100? A)4. 93 A B) 2.00 A C) o.59 A D) 3.05 A E) 11.23 A 4. There are 3 loops in a standing wave in a 4 m string oscillating at 60 Hz What is the wavelength? A) 2.4 m B) 0.75 m C) 1.2m D) 2.67 m E) 6 m 5. A 450 Hz siren is moving at 70 fs toward a person running away from it at 20 fs. What pitch does the person hear? (vsound=1 100 fs) A) 180 Hz B) 1260 Hz C) 471.8 Hz D) 158 Hz E) 900 Hz 6. What is the l. overtone frequency of a 3 ft clarinet? (VM-1 100 fs)--Hz. A) 120 B) 275 C) 467 D) 550 E) 325 7. If a 600 W heater carries a current of 5 A, what is the resistance of the heating element? A) 2.777778 ohm B) 24 ohm C) 20 ohm D) 120 ohm E) 48 ohm 8. An increase of 30 dB increases the sound intensity by a factor of A) 240 B)1080 C) 2700 D) 10000 E) 1000 9. y 6 sin(3t -5x). Velocity of propagation of the wave is A) 3.6 B) 0.6 C) 1.8 D) 3 E) 1.666667 is changing at a rate of 60 A/s? A) 72 mH B) 50 mH C) 480 mH D) 25 mH E) 144 mH 10. What is the self-inductance in a coil which experiences a 3 v induced ent when the current
Answer #1

1) Solution-

Using the equation-

1/f= 1/p+1/q

=> 1/25-1/-80

=36.36cm

Now the di-optical is the inverse of the focal length

D=1/f= 1/0.3636

=>D = 2.75 Diopters